Given: Margin of error (d)= 2.19, Confidence interval (CI) of population mean = 90%, and population standard deviation (σ)= 5 solve for the appropriate sample size.

1
Expert's answer
2022-05-19T18:10:56-0400

The critical value for "\alpha = 0.1" is "z_c = z_{1-\alpha/2} = 1.6449"


"d=z_c\times\dfrac{\sigma}{\sqrt{n}}"

"n=(\dfrac{z_c\times \sigma}{d})^2"

"n=(\dfrac{1.6449\times 5}{2.19})^2"

"n=14"